OpenStudy (anonymous):

Two equations are listed below. Solve each equation and compare the solutions. Choose the statement that is true about both solutions. Equation 1 Equation 2 |5x − 6| = −41 |7x + 13| = 27

OpenStudy (anonymous):

easy he solve it urself

OpenStudy (triciaal):

equation 1 |5x − 6| = −41 -(5x -6) = -41 and 5x -6 = -41 -5x +6 = -41 5x = -41 + 6 -5x = -41 -6 5x = -35 -5x = -47 x = -7 x = 47/5

OpenStudy (triciaal):

equation 2 |7x + 13| = 27 _(7x + 13) = 27 and 7x + 13 = 27 you can continue

OpenStudy (triciaal):

Choose the statement that is true about both solutions.